OpenStudy (irishboy123):

linear system of 1st order DE's

OpenStudy (irishboy123):

say i have a system of linear DE's \(\dot x = x + y\) \(\dot y = x - y\) just made that up so the numbers may be inconvenient; and to be clear, \( \dot x, \dot y\) means \( \frac{dx}{dt}, \frac{dy}{dt}\) ; and \(x = x(t)\), \(y = y(t)\) then \[\left[\begin{matrix}\dot x \\ \dot y\end{matrix}\right] = \left[\begin{matrix}1 & 1 \\ 1 & -1\end{matrix}\right] \left[\begin{matrix}x \\ y\end{matrix}\right]\] with eigenvalues \( \; \pm \sqrt{2}\) but what if it's \(\dot x = x + y \color{red}{+2}\) \(\dot y = x - y \color{red}{+3}\) ie just adding in those constants?? i can't remember but i think i need to use superposition, ie from \[\left[\begin{matrix}\dot x \\ \dot y\end{matrix}\right] = \left[\begin{matrix}1 & 1 \\ 1 & -1\end{matrix}\right] \left[\begin{matrix}x \\ y\end{matrix}\right] + \left[\begin{matrix}2 \\ 3 \end{matrix}\right]\] so separately, i find the particular solution \[\left[\begin{matrix}\dot x \\ \dot y\end{matrix}\right] = \left[\begin{matrix}2 \\ 3 \end{matrix}\right]\] \[\rightarrow \left[\begin{matrix} x \\ y\end{matrix}\right]_p = t \left[\begin{matrix}2 \\ 3 \end{matrix}\right] + \left[\begin{matrix}C_x \\ C_y \end{matrix}\right]\] **i'm looking at finding fixed points; and sketching the phase diagram**. so, for the phase diagram, i then find a way to build that extra "particular" bit in? doesn't feel right though. isn't there a better way to do this?!?! have googled like crazy and nothing there that i can find. many thanks in advance 😋

OpenStudy (anonymous):

Wow Thats quit hard!

ganeshie8 (ganeshie8):

undetermined coefficients should work smoothly in finidng a particular solution : let \(\vec{c}=\begin{pmatrix}c_1\\c_2\end{pmatrix}\) be a particular solution. plugging this into the system gives \[\left[\begin{matrix} 0\\ 0\end{matrix}\right] = \left[\begin{matrix}1 & 1 \\ 1 & -1\end{matrix}\right] \left[\begin{matrix}c_1 \\ c_2\end{matrix}\right] + \left[\begin{matrix}2 \\ 3 \end{matrix}\right] \] we can solve \(\vec{c}\) as the matrix is invertible..

OpenStudy (irishboy123):

thanks ganesh i eventually did this by ID'ing the fixed points \(\dot x = x + y +2 = 0 \\ \dot y = x - y +3 = 0 \) \(\implies x,y = -\frac{5}{2}, \frac{1}{2}\) then re-centering round the fixed point gives \(X = x+\frac{5}{2}, Y = y -\frac{1}{2}; \quad \dot X = \dot x, \; \dot Y = \dot y \) so \(\dot X = X - \frac{5}{2} + Y + \frac{1}{2} +2 = X+Y= 0 \\ \dot Y = X - \frac{5}{2} - (Y + \frac{1}{2} ) +3 = X-Y=0 \) so we have the same eigenvalues..... and hey presto 💥 your solution yields \(c_1 + c_2 = 2\\c_1 - c_2 = 3\) \(c_1, c_2 = \frac{5}{2}, -\frac{1}{2} \) the method i am using is intended to unravel phase diagrams for non linear. i think i have overlooked the linear approach, which is not good ;-( i think i am explaining myself really badly here, but your post was a tremendous help. i will try make this clearer tomorrow when i have more time to sift through it.....

OpenStudy (irishboy123):

i have to close this to open another but i still intend to come back to it......if that's ok.